OpenStudy (anonymous):

\[(x ^{2}- xy+y ^{2})dx-xydy=0\] solve de

OpenStudy (turingtest):

\[(x ^{2}- xy+y ^{2})dx-xydy=0\]\[(x ^{2}- xy+y ^{2})=xy\frac{dy}{dx}=xyy'\]\[\frac xy+1+\frac yx=y'\]now we sub in \(\large v=\frac yx\) from which we get\[y=vx\to y'=v+xv'\]and plug that into the DE

OpenStudy (anonymous):

thank you..i ahve a problem the the DE where u cant saparate xy

OpenStudy (turingtest):

do you see how I did it above? do you get it now or do you need more help?

OpenStudy (anonymous):

ive solve it

OpenStudy (anonymous):

how do i do this \[3x(xy-2)dx+(x ^{3}+2y)dy =0\]

OpenStudy (turingtest):

what is the difference? should be the same procedure

OpenStudy (anonymous):

ok. let me try it

OpenStudy (turingtest):

Man, sorry, this is an exact differential equation, so that previous method won't work this time I have to go teach a class right now, so here's a link for exact DE's http://tutorial.math.lamar.edu/Classes/DE/Exact.aspx you still need to get dy/dx together first though good luck!